# Say My Name: a Model’s Bias Discovery Framework

### v1

@@ 3.2 Bias Naming
Figure 2: Pipeline for SaMyNa. Given a model, we can tell on either $\mathcal{D}^{\text{train}}$or $\mathcal{D}^{\text{val}}$ what are the correct (with green border) and the incorrectly (red border) classified samples. Amongst these, we first perform a sample subset selection looking at the latent space of the model under analysis and choosing through k-medoids the most representative samples for the learned class. Then, we employ a captioner to get a textual description of these samples. Among these descriptions, we identify recurrent keywords and, in parallel, working in the latent space of a text encoder, we extract the mean description for the learned classes, cleansed from common features within the dataset. We finally compare this representation with the embedding of the keywords, identifying learned correlations aside from the target.
We present here SaMyNa, our bias naming approach starting from a trained model $\mathcal{M}$ from which we aim to run our bias naming tool. Fig. 2 proposes an overview of the main pipeline we used, consisting of the following steps:
@@
    Samples captioning. Once the samples are selected, a multimodal LLM captioning tool is used to extract a textual description of each sample (Sec. 3.2.2).
@@
    Keywords selection. Starting from the computed captions, we mine keywords in common from the textual description of the samples within the same learned class (Sec. 3.2.3).
@@
    Learned class embedding. Starting from a textual description of the samples, we can extract the shared information between the correctly classified samples and the incorrectly classified ones: this will constitute the embedding for a potential bias (Sec. 3.2.4).
@@
    Keywords ranking. The embedding of the learned class is compared with the embedding of recurrent keywords in the captions and we get a ranking for the keywords most aligned with the learned class (Sec. 3.2.5).
@@ 3.2.1 Samples subset selection
Given $\mathcal{M}$, for a given target class $c$, we extract the pool of correctly classified samples $\mathcal{D}^{\text{correct}}(c)$ and samples misclassified as $c$ $\mathcal{D}^{\text{misclass}}(c)$.222please note that we have dropped the “train” subscript from the $\mathcal{D}$ partitioning as this pipeline will work equally well also when using a validation set. Provided that $\mathcal{M}$ clusters both $\mathcal{D}^{\text{correct}}(c)$ and $\mathcal{D}^{\text{misclass}}(c)$ together, our hypothesis is that these two share a common set of features, behind which we might find a bias. In the typical deployment scenario, the correctly classified examples are abundant, and, for instance, $\mathcal{M}$ projects them in a very narrow neighborhood of its latent space. We build on top of this observation and run a k-medoid algorithm to reduce the cardinality of correctly classified samples. Our long-range objective will be indeed to capture the set of features that are correctly learned by the model, and k-medoids is a natural choice to have a good coverage of the latent space for $\mathcal{M}$.
@@ 3.2.2 Samples captioning
At this point, we will generate captions from the selected samples. To do this, we use a pre-trained multimodal large language model that takes as input both a prompt and an image. The choice of the prompt is kept generic and asks to generate a textual description of the content of the image, providing some context of the target task. Through all our experiments we will use a large-scale captioner, given that biases might hide in very specific characteristics of the provided images. Images are also preprocessed to ensure they meet the requirements of the model, such as size.
@@ 3.2.3 Keywords selection
From the captions obtained in the previous step (Sec. 3.2.2), we select recurrent keywords. First, we perform some NLP standard processing, including for example stop-words removal. Each caption is word-level tokenized [34], and every token is considered a potential keyword. Then, we count the frequencies of the obtained keywords within the same class $c$. Lastly, we filter out the keywords appearing in the captions of less than the $f_{\text{min}}$ fraction of samples. The remaining keywords are aggregated and constitute a keywords proposal pool $\Psi$.
@@ 3.2.4 Learned class embedding
In parallel to keywords selection, we aim at having a representation for the learned class, disentangled from the specific domain $\mathcal{M}$ is trained to. To do this, we work in the embedding space of a pre-trained text encoder. From the generated captions we obtain, for each class $c$, the embedding matrices ${}{E^{\text{correct}}(c)\in\mathbb{R}^{|\mathcal{D}^{\text{correct}}(c)|\times Z}}$ and ${}{E^{\text{misclass}}(c)\in\mathbb{R}^{|\mathcal{D}^{\text{misclass}}(c)|\times Z}}$, where and $Z$ is the dimensionality of the embedding vectors.
Our goal here is to calculate the embeddings ${}{E^{*}(c)\in\mathbb{R}^{Z}}$ semantically representing the learned representation of the class $c$. Not only: we would like to disentangle this from the features in common to all the classes learned from the model, given that $\mathcal{M}$ could be trained (and tested) to fit a specific domain, which in such specific case it would not constitute a bias but rather a feature. For this, we first calculate the average embedding $E(c)$ for a specific learned class:
$E(c)=\frac{\sum_{i=1}^{|\mathcal{D}^{\text{correct}}(c)|}\sum_{j=1}^{|\mathcal{D}^{\text{misclass}}(c)|}[E_{i}^{\text{correct}}(c)+E_{j}^{\text{misclass}}(c)]}{2\left[|\mathcal{D}^{\text{correct}}(c)|\cdot|\mathcal{D}^{\text{misclass}}(c)|\right]}.$ (3)
$E(c)$ will now contain all the common features of the class $c$. However, it will also contain some information shared in the entire dataset (for example common characteristics of the different classes). From this, we can extract the embedding without the shared information from the dataset through:
@@
The intuition of this approach originates from the arithmetic and semantic properties of natural language latent spaces [35]. To provide a realistic example, consider the task of gender recognition from facial pictures, in which the hair color is a spurious correlation. $E(c)$ might contain features related to concepts such as “blonde” and “face”. As we are only interested in the former, computing $E^{*}(c)$ is an effective solution to filter out the shared information “face”.
@@ 3.2.5 Keywords ranking
where $\psi^{\text{embed}}$ is the embedding of the keyword in the same latent space used to calculate $E^{*}(c)$. This tells us how much the concept is embodied by the proposed keywords. Based on the ranking, we will obtain a set of keywords that correlate with the learned class $c$, and others that become decorrelated as they embody some knowledge shared through all the classes (as filtered in (4)). For this, we introduce a hyper-parameter $t_{\text{sim}}>0$ that thresholds the relevant keywords for the learned class $c$, based on the similarity score. Finally, as post-processing, we filter all the keywords related to the ground-truth target class the model was aiming at learning: the final ranking we obtain embodies the set of features that correlate with the learned class $c$, from which an end user of the system can deduce the presence of a bias.
@@ 4 Empirical Results
We provide here the main results obtained. We highlight that, for visualization purposes, all the figures contain up to the top nine keywords identified by SaMyNa: the full results, along with the ablation study, are presented in the Supplementary material. For our experiments, we have employed an NVIDIA A5000 with 24GB of VRAM, except for the captioning step for which we have employed an NVIDIA A100 equipped with 80GB of VRAM. The source code, in the Supplementary material, will be open-sourced upon acceptance of the article.
@@ 4.1 Setup
Models tested. We tested the most popular architectures benchmarked from the debiasing literature: ResNet-18 for CelebA and BAR, and ResNet-50 for Waterbirds and ImageNet-A. All the models are pre-trained on ImageNet-1K, with architecture and weights provided by torchvision. On ImageNet-A, models are run only in inference, as we are interested in mining biases already existing in the original pre-trained models, while for all the other experiments we apply the training procedure described in Sec. 3.1. For this step, we train with a batch size of 128 and a learning rate of 0.001 for Waterbirds, as done in [24]; for CelebA, we use a batch size of 256 and a learning rate of 0.0001, following [8]. For both, we employ SGD with Nestorov momentum set to 0.9. Finally, for BAR, we employ a batch size of 256 and a learning rate of 0.001, with Adam as optimizer [36].
Captioning. For the captioning model, we used LLaVA-NeXT [37]333 https://huggingface.co/llava-hf/llava-v1.6-34b-hf in its 34B configuration, quantized in 8 bits. The prompts we used to generate the captions are personalized for each dataset and their length is limited to 300 tokens. All the employed captions can be consulted in the Supplementary material.
@@
Class and keywords embedding. For this part, we used the Sentence-BERT model [^4] https://huggingface.co/sentence-transformers/paraphrase-MiniLM-L12-v2 from the sentence-transformers [38] library to generate 384-dimensional embeddings of the captions. The minimum frequency for the keywords $f_{\text{min}}$ is 15%. For the correctly classified samples, we set $k=10$ and $t_{\text{sim}}$ is set to 0.2.
Figure 3: Similarity scores for the Waterbirds dataset.
Figure 4: Similarity scores for the CelebA dataset (trained on the blonde hair attribute).
4.2 Datasets
Before we present the results obtained with our bias naming pipeline, we briefly describe the datasets employed in this study: Waterbirds [24], CelebA [39], BAR [8], and ImageNet-A [40].
@@ 4.2 Datasets
Waterbirds. Waterbirds is an image dataset introduced in [24] to test optimization methods able to be robust against distribution shifts. The associated task is to classify the habitat of bird species, divided into waterbirds and landbirds. In the training set, though, 95% of waterbirds are associated with a water background, and 95% of landbirds are set on land. Only 5% of the two classes’ samples are presented with an opposite background. This results in a potentially strong spurious relation between the target label and the background.
@@
CelebA. CelebA [39] is one of the most popular datasets of face images, depicting celebrity individuals with multiple samples per subject and equipped with extensive annotations regarding roughly 40 attributes, encompassing several face and style characteristics. Regardless of its popularity, it is notoriously affected by several biases [8, 11]. In this work, we analyze the task of classifying whether a person has blond hair or not, for which the attribute female is not uniformly represented and instead is spuriously correlated with the blond class.
@@
BAR. Biased Action Recognition (BAR) is an action recognition dataset crafted by Nam et al. in [8]. The target classes of BAR (Climbing, Diving, Fishing, Racing, Throwing, Vaulting) present a strong bias towards the setting where they are performed. For instance, the large majority of samples from the class Racing is depicted in a circuit track context, while in the test set, we can find many off-road settings on which deep classification models fail to generalize. This dataset does not provide bias group annotations and does not provide a proper validation set.
@@
ImageNet-A. ImageNet-A is a collection of 7,500 real-world images sharing the same category of a 200-class subset of ImageNet, onto which deep models systematically fail to output the correct prediction. Originally introduced in [40] for testing adversarial model robustness, it is also commonly adopted in the context of model bias [13, 41].
4.3 Main results
@@ 4.3.1 Naming the bias at training time
We begin by discussing the results of our bias naming pipeline when applied in a model’s training process on Waterbirds, CelebA, and BAR.
@@
Waterbirds. The barplot in Fig. 3 shows the candidate bias keywords for Waterbirds, alongside their relative similarity value. We can observe how the obtained keywords are mainly related to the background information (tree and forest for landbirds; sea and ocean for waterbirds). Most importantly, the top keywords display high similarity values, indicating a high correlation with their class targets. We deduce that the model suffers from a bias about image backgrounds, which indeed is the case for Waterbirds. It is also worth noticing how the top similarities differ among the two classes. We hypothesize two possible factors causing it: (i) model bias towards sea is stronger than the one towards tree, as it is constituted by simpler visual patterns, easier to learn for the network; (ii) the landbirds class has a much larger population, thus allowing for the bias on this class to be averaged over more instances than the waterbirds case.
@@
CelebA. In analyzing the possible biases of our vanilla model on CelebA (see Fig. 4), we find that the top-1 keyword for both classes represents a gender: male/man for class not blonde (with similarity $\approx 0.4$), and woman for class blonde (similarity of $0.49$). Additionally, we find among the blonde class, keyword terms typically associated with gender stereotypes, such as makeup and lipstick, with moderately high similarity. This suggests that the vanilla model is not only biased regarding the classification task but also incorporates unfair features, arising from societal biases reflected in the training data.
@@
BAR. Bar presents a more challenging situation, due to the complete absence of bias group annotations. Regardless, the output of our approach still provides insights about the biases captured by the model. In the training class climbing, scenes where the subject is ascending on rocks, are overly represented, and thus the vanilla model has wrongfully learned to rely on the presence of rocky backgrounds. This is reflected by the top three keywords in Fig. 5 (cliff, rock and rocks), all having similarities above $0.4$. Similar considerations can be made on the other classes: pitch, batter, glove suggest an even stronger bias for the class throwing towards a specific sport, baseball, which is also the second most correlated keyword. The same can be said for scuba in the diving class. Keywords for other classes show slightly lower maxima. Still, we can measure relevant correlations with the presence of cars and circuit races for racing, as well as boats and water (e.g. rivers, sea) for fishing.

### v2

@@ 3.2 Bias Naming
We present here SaMyNa, our bias naming approach. Fig. 2 proposes an overview of the main pipeline we used, consisting of the following steps:
@@
    Samples captioning. We then generate textual descriptions for the selected samples (Sec. 3.2.2).
@@
    Keywords selection. Starting from the captions, we construct a pool of keywords by applying a lightweight word frequency filter. Heavy filtering is done as a last step to better account for synonyms (Sec. 3.2.3).
@@
    Learned class embedding. Simultaneously, we generate embeddings for the captions. We use these embeddings, along with the corresponding target and predicted classes, to calculate learned class embeddings that represent the biases of $\mathcal{M}$ if present (Sec. 3.2.4).
@@
    Keywords ranking. We then embed the keywords and rank them according to their similarity with the learned class embeddings. Finally, low-ranking keywords can be filtered out (Sec. 3.2.5).
@@ 3.2.1 Samples Subset Selection
Given $\mathcal{M}$, for a given target class $c$, we extract the pool of correctly classified samples $\mathcal{D}^{\text{correct}}(c)$ and the pool of samples misclassified as class $c$, which we call $\mathcal{D}^{\text{misclass}}(c)$.111please note that we have dropped the “train” subscript from the $\mathcal{D}$ partitioning as this pipeline will work equally well also when using a validation set. Provided that $\mathcal{M}$ clusters both $\mathcal{D}^{\text{correct}}(c)$ and $\mathcal{D}^{\text{misclass}}(c)$ together, our hypothesis is that these two share a common set of features, behind which we might find a bias. In the typical deployment scenario, the correctly classified examples are abundant, and, for instance, $\mathcal{M}$ may project them in a very narrow neighborhood of its latent space. We build on top of this observation, and we cluster correctly classified samples using the $k$-medoid algorithm to reduce the cardinality of correctly classified samples while maintaining a good coverage of the model’s latent space. The choice of using $k$-medoid is natural, since it provides samples as output instead of centroids (which may be arbitrary points in feature space). As our long-range objective is to capture the set of features learned by the model to make its predictions, we also select a sample of examples misclassified as class $c$ by sorting them according to $d_{t^{*},n}$, selecting the $k$ ones that were misclassified with more confidence. We refer to the $k$ samples resulting from running $k$-medoid on the correctly classified samples as $\mathcal{S}^{\text{correct}}(c)$, while the pool of $k$ samples misclassified as $c$ is denoted as $\mathcal{S}^{\text{misclass}}(c)$.
@@ 3.2.2 Samples Captioning
At this point, we will generate captions from the selected samples. To do this, we use a pre-trained multimodal large language model that takes as input both a prompt and an image. We used the following generic captioning prompt for all the experiments: “Describe this image in detail. Pay particular attention to the subject and where they are. Describe as many details as possible”.
We decided to employ a large-scale image captioner in all our experiments, given that biases might hide in the subtle characteristics of the provided images (see Sec. B of the Appendix).
@@ 3.2.3 Keywords Selection
From the captions obtained in the previous step (Sec. 3.2.2), we select candidate keywords. First, we perform some NLP standard processing, consisting of lower-case conversion, word-level tokenization (Bird et al., 2009), and stop-words removal. Then, for each token $w$ and class $c$, we count the number $f_{w}(c)$ of captions of class $c$ that contain $w$ at least once. Lastly, given $f_{\text{min}}$ between 0 and 1, we filter out the token $w$ if $\text{max}_{c}[f_{w}(c)]$ is less than $f_{\text{min}}\cdot[|\mathcal{S}^{\text{correct}}(c)|+|\mathcal{S}^{\text{misclass}}(c)|]$. The remaining tokens are aggregated into a keywords proposal pool $\Psi$.
@@ 3.2.4 Learned Class Embedding
In parallel to keywords selection, we aim at having a representation for the learned class, disentangled from the specific domain $\mathcal{M}$, which is trained to. To do this, we work in the embedding space of a pre-trained text encoder. From the generated captions we obtain, for each class $c$, the embedding matrices $E^{\text{correct}}(c)\in\mathbb{R}^{|\mathcal{S}^{\text{correct}}(c)|\times Z}$} and $E^{\text{misclass}}(c)\in\mathbb{R}^{|\mathcal{S}^{\text{misclass}}(c)|\times Z}$}, where $Z$ is the dimensionality of the embedding vectors. These matrices are simply the results of stacking the embedding vectors coming from the text encoder applied to the various captions of the samples in $\mathcal{S}^{\text{correct}}(c)$ and $\mathcal{S}^{\text{misclass}}(c)$. Our goal here is to calculate the embeddings $E^{*}(c)\in\mathbb{R}^{Z}$} semantically representing the learned representation of the class $c$. Not only that, we would like to disentangle this from the features in common to all the classes learned from the model, given that $\mathcal{M}$ could be trained (and tested) to fit a specific domain, which in such a specific case would not constitute a bias but rather a feature. For this, we first calculate the embedding $E(c)$ for a specific learned class:
$E(c)=\frac{\sum_{i=1}^{|\mathcal{S}^{\text{correct}}(c)|}\sum_{j=1}^{|\mathcal{S}^{\text{misclass}}(c)|}[E_{i}^{\text{correct}}(c)+E_{j}^{\text{misclass}}(c)]}{2\left[|\mathcal{S}^{\text{correct}}(c)|\cdot|\mathcal{S}^{\text{misclass}}(c)|\right]}.$ (3)
This equation computes the average of the means across all pairs of embeddings, where each pair consists of a correctly classified example of class $c$ and an example that was misclassified as class $c$. We adopt this pairwise formulation instead of a simple average to improve robustness in scenarios with significant imbalance between correctly and incorrectly classified samples. This imbalance usually happens when $k$ is greater than the number of misclassified examples for a given class. Additionally, because each pair includes embeddings from examples with different ground-truth classes, this approach helps discourage the encoding of the target class into the learned class embedding. $E(c)$ will now contain all the common features of the class $c$. However, it will also contain some information shared in the entire dataset (for example common characteristics of the different classes). From this, we can extract the embedding without the shared information from the dataset through:
@@
The intuition of this approach originates from the arithmetic and semantic properties of natural language latent spaces (Mikolov et al., 2013). To provide a realistic example, consider the task of gender recognition from facial pictures, in which the hair color is a spurious correlation. $E(c)$ might contain features related to concepts such as “blonde” and “face”. As we are only interested in the former, computing $E^{*}(c)$ is an effective solution to filter out the shared information “face”. We demonstrate the effectiveness of equation 4 through an ablation study in Sec. C.2.1 of the Appendix.
@@ 3.2.5 Keywords Ranking
where $\psi^{\text{embed}}$ is the embedding of the keyword in the same latent space used to calculate $E^{*}(c)$. This tells us how much the concept is embodied by the proposed keywords. Based on the ranking, we will obtain a set of keywords that correlate with the learned class $c$, and others that become decorrelated as they embody some knowledge shared through all the classes (as filtered in equation 4). For this, we introduce a hyper-parameter $t_{\text{sim}}>0$ that thresholds the relevant keywords for the learned class $c$, based on the similarity score. The final ranking we obtain embodies the set of features that correlate with the learned class $c$, from which an end user of the system can deduce the presence of a bias. In some cases, the ranking may also contain keywords related to the target class, but such keywords are easily ignored by the end user.
@@ 4 Empirical Results
We provide here the main results obtained. We highlight that, for visualization purposes, all the figures contain up to the top nine keywords identified by SaMyNa: additional ablation studies and the full results are presented in Sec. C and G of the Appendix, respectively. For our experiments, we have employed an NVIDIA A5000 with 24GB of VRAM, except for the captioning step for which we have employed an NVIDIA A100 equipped with 80GB of VRAM. The source code, attached to the submission, will be open-sourced upon acceptance of the article.
@@ 4.1 Setup
Models tested. We tested the most popular architectures benchmarked from the debiasing literature: ResNet-18 for CelebA and BAR, and ResNet-50 for Waterbirds and ImageNet-A. All the models are pre-trained on ImageNet-1K, with architecture and weights provided by torchvision. On ImageNet-A, models are run only in inference, as we are interested in mining biases already existing in the original pre-trained models, while for all the other experiments we apply the training procedure described in Sec. 3.1. For this step, we train with a batch size of 128 and a learning rate of 0.001 for Waterbirds, as done in (Sagawa* et al., 2020); for CelebA, we use a batch size of 256 and a learning rate of 0.0001, following (Nam et al., 2020). For both, we employ SGD with Nesterov, set to 0.9. Finally, for BAR, we employ a batch size of 256 and a learning rate of 0.001, with Adam as optimizer (Kim et al., 2021).
Captioning. For the captioning model, we used LLaVA-NeXT (Liu et al., 2024) [^2] https://huggingface.co/llava-hf/llava-v1.6-34b-hf in its 34B configuration, quantized in 8 bits. Before feeding our input images to the image captioner, we apply the corresponding preprocessing transform provided by the huggingface library.
@@
Class and keywords embedding. For this part, we used a MiniLM model [^3] https://huggingface.co/sentence-transformers/paraphrase-MiniLM-L12-v2 from the sentence-transformers (Reimers & Gurevych, 2019) library to generate 384-dimensional embeddings of the captions. The minimum frequency for the keywords $f_{\text{min}}$ is 15%. For the correctly classified samples, we set $k=10$ and $t_{\text{sim}}$ is 0.2.
Datasets. For our study, we employ the following datasets: Waterbirds (Sagawa* et al., 2020), CelebA (Liu et al., 2015), BAR (Nam et al., 2020), and ImageNet-A (Hendrycks et al., 2021).
@@
Waterbirds is an image dataset introduced in Sagawa* et al. (2020) to test the robustness of optimization methods against distribution shifts. The associated task is to classify the habitat of bird species, divided into waterbirds and landbirds. In the training set, though, 95% of waterbirds are associated with a water background, and 95% of landbirds are set on land. Only 5% of the two classes’ samples are presented with an opposite background. This results in a potentially strong spurious relation between the target label and the background.
@@
CelebA Liu et al. (2015) is one of the most popular datasets of face images, depicting celebrity individuals with multiple samples per subject and equipped with extensive annotations regarding roughly 40 attributes, encompassing several face and style characteristics. Regardless of its popularity, it is notoriously affected by several biases Nam et al. (2020); Barbano et al. (2023). In this work, we analyze the task of classifying whether a person has blond hair or not, for which the attribute female is not uniformly represented, but spuriously correlated with the blond class. A deep network trained with standard techniques exhibits a strong bias in predicting subjects with female perceived gender as blond, while male appearing subjects are often classified as not blond.
@@
Biased Action Recognition (BAR) is an action recognition dataset crafted by Nam et al. in Nam et al. (2020). The target classes of BAR (Climbing, Diving, Fishing, Racing, Throwing, Vaulting) present a strong bias towards the setting where they are performed. For instance, the large majority of samples from the class Racing is depicted in a circuit track context, while in the test set, we can find many off-road settings on which deep classification models fail to generalize. This dataset does not provide bias group annotations and does not provide a proper validation set, making it impossible to compute a Worst Group accuracy.
@@
ImageNet-A is a collection of 7,500 real-world images sharing the same category of a 200-class subset of ImageNet, onto which deep models systematically fail to output the correct prediction. Originally introduced in Hendrycks et al. (2021) for testing adversarial model robustness, it is also commonly adopted in the context of model bias Bahng et al. (2020b); Kim et al. (2022b).
Figure 3: Similarity scores for Waterbirds (a) and CelebA (b).
4.2 Bias Discovery
Our empirical validation on bias discovery is here divided into naming the bias during training (Sec. 4.2.1) and naming it post-hoc, at inference (Sec. 4.2.2). In Sec. 4.3 we provide a description of how our keywords can be used to perform bias mitigation, and in Sec. 5.2 we show how SaMyNa behaves in an unbiased case.
@@ 4.2.1 Naming the Bias at Training Time
We begin by discussing the results of SaMyNa when applied in a model’s training on Waterbirds, CelebA, and BAR.
@@
Waterbirds. The barplot in Figure˜3(a) shows the candidate bias keywords for Waterbirds, alongside their relative similarity value. We can observe how the obtained keywords are mainly related to the background information (forest and foliage for landbirds; ocean and beach for waterbirds). Most importantly, the top keywords display high similarity values, indicating a high correlation with their class targets. We deduce that the model suffers from a bias about image backgrounds, which indeed is the case for Waterbirds. It is also worth noticing how the top similarities differ among the two classes. We hypothesize two possible factors causing it: (i) model bias towards ocean is stronger than the one towards forest, as it is constituted by simpler visual patterns, easier to learn for the network; (ii) the landbirds class has a much larger population, thus allowing for the bias on this class to be averaged over more instances than the waterbirds case.
@@
CelebA. In analyzing the possible biases of our vanilla model on CelebA (see Figure˜3(b)), we find that the top-1 keyword for both classes represents a gender: male/man for class not blonde (with similarity $\approx 0.4$), and woman for class blonde (similarity of $\approx 0.4$). Additionally, we find among the blonde class, keyword terms typically associated with gender stereotypes, such as mascara, makeup, and lipstick, with moderately high similarity. This suggests that the vanilla model is not only biased regarding the classification task but also incorporates unfair features, arising from societal biases reflected in the training data.
@@
BAR. Bar presents a more challenging situation, due to the complete absence of bias group annotations. Regardless, the output of our approach still provides insights about the biases captured by the model. In the training class climbing, scenes where the subject is ascending on rocks are overly represented, and thus the vanilla model has wrongfully learned to rely on the presence of rocky backgrounds. This is reflected by the top three keywords not related to the target class in Figure˜4 (cliff, rock and steep), all having high similarities. Similar considerations can be made on the other classes: pitch, baseball, pitcher, and batter suggest an even stronger bias for the class throwing towards a specific sport, baseball, which is also the second most correlated keyword. The same can be said for scuba and underwater in the diving class. Keywords for other classes show slightly lower maxima. Still, we can measure relevant correlations with the presence of cars and circuit races for racing, as well as boats and water (e.g. rivers, ocean, lake) for fishing. Complete outputs of the keyword extraction process are available in the Appendix.
Method No Val.Set
in Bias-Id
Unsup. Bias
Named
CelebA (Hair Color) Waterbirds BAR
Average Worst Group Average Worst Group Average
LISA Yao et al. (2022) $92.40$ $89.30$ $91.80$ $89.20$
GroupDRO Sagawa* et al. (2020) $92.90\pm 0.20$ $88.90\pm 2.30$ $93.50\pm 0.30$ $91.40\pm 1.10$
George Sohoni et al. (2020) $94.60$ $54.90\pm 1.90$ $95.70$ $76.20\pm 2.00$
JTT Liu et al. (2021) $88.10$ $81.50\pm 1.70$ $89.30$ $83.80\pm 1.20$ $68.53\pm 3.29$
CNC Zhang et al. (2022b) $89.90$ $88.80\pm 0.90$ $90.90$ $88.50\pm 0.30$
B2T+GDRO Kim et al. (2024) $93.20$ $90.40\pm 0.90$ $91.50$ $\bf 90.70\pm 0.30$ $68.54\pm 0.86^{*}$
ERM $\bf 94.90$ $47.70\pm 2.10$ $\bf 97.30$ $62.60\pm 0.30$ $51.85\pm 5.92$
LfF Nam et al. (2020) $84.24$ $81.24\pm 1.38$ $91.20$ $78.00$ $62.98\pm 2.76$
DebiAN Li et al. (2022) $84.00$ $52.90\pm 4.70$ $69.88\pm 2.92$
SaMyNa (ours) + GDRO $92.20\pm 0.01$ $\bf 90.60\pm 0.08$ $91.11\pm 0.44$ $\bf 90.62\pm 0.10$ $\bf 71.26\pm 1.46$
Table 1: Performance of GDRO Debiasing on top of our training-time bias naming pipeline. Column Unsup. indicates if the method uses ground truth bias information. Best results for unsupervised debiasing methods are highlighted in bold. No Val.Set in Bias-Id highlights if the method does not rely on a validation set for inferring subgroups or bias-attributes (green tick, ) or it assumes having one (red cross, ). Bias Named indicates if the method extracts semantic names of found bias attributes. BAR lacks worst-group accuracy due to the absence of bias annotations. Average refers to the unbalanced test accuracy. These results on B2T+GDRO were computed by us.